# DAWN: Dynamic Frame Avatar with Non-autoregressive Diffusion Framework for talking head Video Generation

### v1

@@ Abstract
Talking head generation intends to produce vivid and realistic talking head videos from a single portrait and speech audio clip. Although significant progress has been made in diffusion-based talking head generation, almost all methods rely on autoregressive strategies, which suffer from limited context utilization beyond the current generation step, error accumulation, and slower generation speed. To address these challenges, we present DAWN (Dynamic frame Avatar With Non-autoregressive diffusion), a framework that enables all-at-once generation of dynamic-length video sequences. Specifically, it consists of two main components: (1) audio-driven holistic facial dynamics generation in the latent motion space, and (2) audio-driven head pose and blink generation. Extensive experiments demonstrate that our method generates authentic and vivid videos with precise lip motions, and natural pose/blink movements. Additionally, with a high generation speed, DAWN possesses strong extrapolation capabilities, ensuring the stable production of high-quality long videos. These results highlight the considerable promise and potential impact of DAWN in the field of talking head video generation. Furthermore, we hope that DAWN sparks further exploration of non-autoregressive approaches in diffusion models. Our code will be publicly at https://github.com/Hanbo-Cheng/DAWN-pytorch.
@@ 1 Introduction
Recently, Diffusion Models (DM) (Ho et al., 2020) have achieved significant success in video and image generation tasks (Rombach et al., 2022; Ho et al., 2022b; a; Peebles & Xie, 2023; Ni et al., 2023). However, their application in talking head generation (Shen et al., 2023; Bigioi et al., 2024) still faces several challenges. Although many methods yield high-quality results, most of them rely on autoregressive (AR) (Tian et al., 2024; Ma et al., 2023) or semi-autoregressive (SAR) (Xu et al., 2024b; He et al., 2023) strategies. In each iteration, AR generates one frame, while SAR generates a fixed-length video segment. The two strategies significantly slow down the inference speed and fail to adequately utilize contextual information from future frames, which leads to constrained performance and potential error accumulation, especially in long video sequences (Stypułkowski et al., 2024).
@@ 2 Related works
Audio-driven talking head generation. Initial approaches for talking head generation employed deterministic models to map audio to video streams (Fan et al., 2016), with later methods introducing generative models such as GANs (Isola et al., 2016), VAEs (Kingma & Welling, 2022), and diffusion models (DMs) (Ho et al., 2020). GAN-related methods (Vougioukas et al., 2020; Pumarola et al., 2018; Hong et al., 2022) improved visual realism but faced convergence and mode collapse issues (Xia et al., 2022). Following this, VAEs (Kingma & Welling, 2022) generated 3D priors like 3D Morphable Models (3DMM) (Blanz & Vetter, 1999) followed by high-fidelity rendering (Ren et al., 2021), which limited the realism and vividness. In contrast, DMs have been introduced into talking head generation due to their good convergence, excellent generation performance, and diversity. Stypułkowski et al. (2024) presented a DM-based talking head generation solution using an AR strategy to generate videos frame-by-frame iteratively. Subsequently, Tian et al. (2024) improved this AR strategy by incorporating motion conditions extracted by a VAE-based network as priors for each iteration, effectively mitigating degradation issues. Concurrently, Xu et al. (2024b); He et al. (2023) advocated for motion modeling instead of image modeling. They utilized DMs to iteratively generate latent motion representations over a fixed number of frames in a SAR manner, subsequently converting these motion representations into video frames. While most diffusion-based methods produce promising results, their AR or SAR strategies incur slow generation speeds and collapse in long-video generation. Although methods like Tian et al. (2024); Xu et al. (2024b) alleviate issues such as inconsistencies in content across iterations and long video generation collapse, the risk of error accumulation remains unresolved. While methods like Du et al. (2023) use identity-specific NAR strategies, to the best of our knowledge, none have addressed NAR talking head generation for arbitrary identities. Consequently, we propose a novel NAR dynamic frame generation framework based on DM, which aims to achieve the low-cost and high-quality rapid generation of realistic talking head video through a clip of audio and arbitrary portrait.
@@ 3.3 Conditional Audio2Video Flow Diffusion Model
Through the LFG in Section 3.2, we can specify a ${\bm{x}}_{\text{src}}$, and extract the identity-agnostic motion representations ${\bm{z}}^{m}_{1:N}$ of the talking head video clip ${\bm{x}}_{1:N}$ as well as the latent code ${\bm{z}}_{\text{src}}$. Therefore, we design the A2V-FDM to generate the motion representations $\hat{{\bm{z}}}^{m}_{1:N}=[\hat{{\bm{f}}}_{1:N},\hat{{\bm{m}}}_{1:N}]$ of each frame relative to ${\bm{x}}_{\text{src}}$:
$\hat{{\bm{z}}}^{m}_{1:N}=\text{DM}({\bm{x}}_{\text{src}},{\bm{y}}_{1:N},\bm{\rho}_{1:N})$ (6)
where the DM refers the diffusion model. After generating $\hat{{\bm{z}}}^{m}_{1:N}$, we use the decoder $\mathcal{D}$ in LFG to reconstruct the $\hat{{\bm{z}}}^{m}_{1:N}$ to the target video frames, via the Equation 4. The structure of A2V-FDM is illustrated in Figure 1 (c). The A2V-FDM model includes a 3D U-Net denoising backbone (Ho et al., 2022b). The residue block in the 3D U-Net contains temporal attention and spatial attention modules, which handle frame-level and pixel-level dependencies respectively. The parameters of the temporal attention module are independent of the input length, so we believe that 3D U-Net can theoretically process video sequences of any length. However, to ease training difficulty, we train on short sequences and aim for long-sequence inference. To enhance the 3D U-Net’s extrapolation ability when handling long sequences, we used Rotary Positional Encoding (RoPE) (Su et al., 2023) instead of traditional absolute position embeddings in the temporal attention module.
@@ 3.4 Audio-Driven Pose and Blink Generation
To enhance the model’s extrapolation capability, we use RoPE as the positional encoding in the decoder, consistent with A2V-FDM. During training, we apply an MSE-based reconstruction loss $L_{\text{rec}}$ and an adversarial loss $L_{\text{GAN}}$ to guide the model in completing basic reconstruction tasks (Isola et al., 2016; Ginosar et al., 2019). Additionally, we employ a KL divergence loss $L_{\text{KL}}$ to ensure that the latent code ${\bm{h}}$ closely approximates a standard Gaussian distribution.
@@
$\hat{{\bm{x}}}_{1:N}=\mathcal{D}(\text{DM}({\bm{x}}_{\text{src}},{\bm{y}}_{1:N},\bm{\rho}_{1:N}))$ (10)
@@ 3.6 Inference
Our inference process consists of four steps: 1) Extract the audio embedding ${\bm{a}}_{1:N}$ from the input audio. 2) Provide the initial pose/blink state $\bm{\rho}_{\text{src}}$, extracted from the source image ${\bm{x}}_{\text{src}}$, to the PBNet. Along with ${\bm{a}}_{1:N}$ and a latent space vector ${\bm{h}}_{1:N}$ sampled from a standard Gaussian distribution, PBNet generates the pose/blink sequences $\hat{\bm{\rho}}_{1:N}$. 3) Input ${\bm{x}}_{\text{src}}$, ${\bm{a}}_{1:N}$, and $\hat{\bm{\rho}}_{1:N}$ into the A2V-FDM, which generates the motion representation sequences $\hat{{\bm{z}}}^{m}_{1:N}$. 4) Finally, decode the video sequence $\hat{{\bm{x}}}_{1:N}$ from ${\bm{x}}_{\text{src}}$ and $\hat{{\bm{z}}}^{m}_{1:N}$.
@@ 4.1 Setup
Implementation details. The architecture of the encoder and decoder in our model aligns with the design proposed by Johnson et al. (2016), while the flow predictor is implemented based on MRAA (Siarohin et al., 2021). The PBNet model is trained using pose and blink movement sequences of 200 frames. During the inference phase of the PBNet model, a local attention mechanism with a window size of 400 is employed. The motion representation space of the A2V-FDM model is $32\times 32\times 3$. The training of the A2V-FDM model is conducted using the TCL strategy, which involves two stages: 1) the model is trained using video clips of 20 frames, with the source image set as the first frame in the clip; 2) training is performed on sequences ranging from 30 to 40 frames. For the inference phase of the A2V-FDM model, local attention with a window size of 80 is applied. In our evaluation, the length of one-time inference for CREMA is dynamic and depends on the ground truth, while for HDTF, it is fixed at 200 frames for better comparison.
Evaluation metrics. We evaluate the performance of our method using various metrics. Specifically, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) to assess the image quality. We utilize the $\text{FVD}_{16}$ and $\text{FVD}_{32}$ scores, which calculate the Fréchet Video Distance (FVD) based on window sizes of 16 and 32 frames, respectively, to evaluate video quality across different temporal scales. Furthermore, we assess the perception loss of lip shape using the confidence score ($\text{LSE}_{\text{C}}$) and distance score ($\text{LSE}_{\text{D}}$) (Chung & Zisserman, 2017). To evaluate the preservation of speaker identity during the generation, we use ArcFace (Deng et al., 2019) to extract features from both the ground truth image and the generated image, and use the cosine similarity (CSIM) between the two as the evaluation metric. Moreover, we employ the Beat Align Score (BAS) (Siyao et al., 2022) to evaluate the synchronization of head motion and audio, and calculate the number of blinks per second (blink/s) to assess the liveliness of the eyes.
@@ 4.2 Overall Comparison
We compared our method with several state-of-the-art methods: Wav2Lip (Prajwal et al., 2020), MakeItTalk (Zhou et al., 2020), Audio2Head (Wang et al., 2021), SadTalker (Zhang et al., 2023), and Diffused Heads (Stypułkowski et al., 2024). To ensure a fair comparison, we provided the same conditions for each baseline and generated the same number of frames in a single generation. The quantitative experimental results are presented in Table 1.
Table 1: Quantitative comparison with several state-of-the-art methods methods on HDTF (Zhang et al., 2021) and CREMA (Cao et al., 2014) datasets. * Wav2Lip generated videos that only contain lip motions, while the rest remain still images. $\uparrow$” indicates better performance with higher values, while $\downarrow$” indicates better performance with lower values. For both BAS and Blink/s, we consider performance to be better when they are closer to the ground truth.
@@
Table 2: Comparison with other generation strategies. The semi-autoregressive (SAR) generation strategy is similar to He et al. (2023). The two temporal resolution (TTR) generation method is mentioned in Harvey et al. (2022).
@@
SAR 11.42 13.00 120.33 210.52 4.34 8.29
TTR 19.25 9.77 95.42 137.14 4.87 8.68
Ours 7.32 9.60 60.34 95.64 6.71 7.94
Table 3: The experiment of extrapolation evaluation. “Inference length" refers to the number of frames generated in a single inference process.
@@
Table 4: Ablation study on TCL and PBNet. The “GT PB" refers to whether to use ground truth pose/blink signal.
@@
According to the results, our method achieves the best performance in FID, $\text{FVD}_{32}$, $\text{FVD}_{16}$, CSIM, BAS, and Blink/s metrics for both the CREMA and HDTF datasets. Additionally, our method achieves nearly best scores for $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$, closely matching those of the Audio2Head and Wav2Lip baselines. These results demonstrate that our method produces talking head videos with the highest visual quality, rhythmic head poses, and natural blink movements. Additionally, it preserves the identity features of the source image to the greatest extent and achieves highly synchronized lip motions with the audio. Additionally, we test the generation speed of the aforementioned methods in Appendix A.2.2, suggesting our method achieves the fastest or near-fastest generation speed and requires significantly less time compared to the previous diffusion-based method, Diffused Heads.
For the qualitative experiment, as shown in Figure 2, we visualize the generation results for each baseline on different datasets. Our method evidently achieves the visual quality most similar to the ground truth, showcasing the most realistic and vivid visual effects. Compared to our method, SadTalker relies on the 3DMM prior, which limits its animation capability to the facial region only, resulting in significant artifacts when merging the face with the static torso below the neck. Additionally, SadTalker exhibits unnatural head pose movements and gaze direction, partially due to limited temporal modeling ability. Wav2Lip only drives the mouth region and cannot generate head poses or blinks. The Audio2Head fails to preserve the speaker’s identity during generation. For the HDTF dataset, the Diffused Heads method collapsed due to the error accumulation.
@@
We compared our non-autoregressive generation strategy with two regular video generation strategies: 1) semi-autoregressive (SAR) generation similar to He et al. (2023), and 2) two-temporal resolution (TTR), which trains two models with different temporal resolutions (Harvey et al., 2022). The time cost represents the time required to generate an 8-second talking head video. The models were evaluated on the CREMA dataset, and the results are shown in Table 2. According to the results, our non-autoregressive method produces videos with the highest quality and fastest speed. The generation speed of the SAR method is faster than TTR, while TTR produces higher video quality than SAR. This is because autoregressive methods suffer from degradation issues during iterative generation. Although TTR somewhat alleviates the degradation issue, it compromises generation speed. In summary, our non-autoregressive method addresses the degradation problem while preserving fast generation speed.

### v3

@@ Abstract
Talking head generation intends to produce vivid and realistic talking head videos from a single portrait and speech audio clip. Although significant progress has been made in diffusion-based talking head generation, almost all methods rely on autoregressive strategies, which suffer from limited context utilization beyond the current generation step, error accumulation, and slower generation speed. To address these challenges, we present DAWN (Dynamic frame Avatar With Non-autoregressive diffusion), a framework that enables all-at-once generation of dynamic-length video sequences. Specifically, it consists of two main components: (1) audio-driven holistic facial dynamics generation in the latent motion space, and (2) audio-driven head pose and blink generation. Extensive experiments demonstrate that our method generates authentic and vivid videos with precise lip motions, and natural pose/blink movements. Additionally, with a high generation speed, DAWN possesses strong extrapolation capabilities, ensuring the stable production of high-quality long videos. These results highlight the considerable promise and potential impact of DAWN in the field of talking head video generation. Furthermore, we hope that DAWN sparks further exploration of non-autoregressive approaches in diffusion models. Our code will be publicly available at https://github.com/Hanbo-Cheng/DAWN-pytorch.
@@ 1 Introduction
Recently, Diffusion Models (DM) (Ho et al., 2020) have achieved significant success in video and image generation tasks (Rombach et al., 2022; Ho et al., 2022b; a; Peebles & Xie, 2023; Ni et al., 2023). However, their application in talking head generation (Shen et al., 2023; Bigioi et al., 2024) still faces several challenges. Although many methods yield high-quality results, most of them rely on autoregressive (AR) (Tian et al., 2024; Ma et al., 2023) or semi-autoregressive (SAR) (Xu et al., 2024c; He et al., 2023) strategies. In each iteration, AR generates one frame, while SAR generates a fixed-length video segment. The two strategies significantly slow down the inference speed and fail to adequately utilize contextual information from future frames, which leads to constrained performance and potential error accumulation, especially in long video sequences (Stypułkowski et al., 2024; Tian et al., 2024; Xu et al., 2024b).
@@ 2 Related works
Audio-driven talking head generation. Initial approaches for talking head generation employed deterministic models to map audio to video streams (Fan et al., 2016), with later methods introducing generative models such as GANs (Isola et al., 2016a), VAEs (Kingma & Welling, 2022), and diffusion models (DMs) (Ho et al., 2020). GAN-related methods (Vougioukas et al., 2020; Pumarola et al., 2018; Hong et al., 2022) improved visual realism but faced convergence and mode collapse issues (Xia et al., 2022). Following this, VAEs (Kingma & Welling, 2022) generated 3D priors like 3D Morphable Models (3DMM) (Blanz & Vetter, 1999) followed by high-fidelity rendering (Ren et al., 2021), which limited the realism and vividness. In contrast, DMs have been introduced into talking head generation due to their good convergence, excellent generation performance, and diversity. Stypułkowski et al. (2024) presented a DM-based talking head generation solution using an AR strategy to generate videos frame-by-frame iteratively. Subsequently, Tian et al. (2024) improved this AR strategy by incorporating motion conditions extracted by a VAE-based network as priors for each iteration, effectively mitigating degradation issues. Concurrently, Xu et al. (2024c); He et al. (2023) advocated for motion modeling instead of image modeling. They utilized DMs to iteratively generate latent motion representations over a fixed number of frames in a SAR manner, subsequently converting these motion representations into video frames. While most diffusion-based methods produce promising results, their AR or SAR strategies incur slow generation speeds and collapse in long-video generation. Although methods like Tian et al. (2024); Xu et al. (2024c) alleviate issues such as inconsistencies in content across iterations and long video generation collapse, the risk of error accumulation remains unresolved. While methods like Du et al. (2023) use identity-specific NAR strategies, to the best of our knowledge, none have addressed NAR talking head generation for arbitrary identities. Consequently, we propose a novel NAR dynamic frame generation framework based on DM, which aims to achieve the low-cost and high-quality rapid generation of realistic talking head video through a clip of audio and arbitrary portrait.
@@ 3.3 Conditional Audio2Video Flow Diffusion Model
Through the LFG in Section 3.2, we can specify a ${\bm{x}}_{\text{src}}$, and extract the identity-agnostic motion representations ${\bm{z}}^{m}_{1:N}$ of the talking head video clip ${\bm{x}}_{1:N}$ as well as the latent code ${\bm{z}}_{\text{src}}$ extracted by $\mathcal{E}$ from ${\bm{x}}_{\text{src}}$. Therefore, we design the A2V-FDM to generate the motion representations $\hat{{\bm{z}}}^{m}_{1:N}=[\hat{{\bm{f}}}_{1:N},\hat{{\bm{m}}}_{1:N}]$ of each frame relative to ${\bm{x}}_{\text{src}}$:
$\hat{{\bm{z}}}^{m}_{1:N}=\text{DM}(\mathcal{E}({\bm{x}}_{\text{src}}),{\bm{y}}_{1:N},\bm{\rho}_{1:N})$ (6)
where the DM refers the diffusion model and $\bm{\rho}_{1:N}$ is the pose/blink signal. After generating $\hat{{\bm{z}}}^{m}_{1:N}$, we use the decoder $\mathcal{D}$ in LFG to reconstruct the $\hat{{\bm{z}}}^{m}_{1:N}$ to the target video frames, via the Equation 4. The structure of A2V-FDM is illustrated in Figure 1 (c). The A2V-FDM model includes a 3D U-Net denoising backbone (Ho et al., 2022b). The residue block in the 3D U-Net contains temporal attention and spatial attention modules, which handle frame-level and pixel-level dependencies respectively. The parameters of the temporal attention module are independent of the input length, so we believe that 3D U-Net can theoretically process video sequences of any length. However, to ease training difficulty, we train on short sequences and aim for long-sequence inference. To enhance the 3D U-Net’s extrapolation ability when handling long sequences, we used Rotary Positional Encoding (RoPE) (Su et al., 2023) instead of traditional absolute position embeddings in the temporal attention module.
@@ 3.4 Audio-Driven Pose and Blink Generation
To enhance the model’s extrapolation capability, we use RoPE as the positional encoding in the decoder, consistent with A2V-FDM. During training, we apply an MSE-based reconstruction loss $L_{\text{rec}}$ and an adversarial loss $L_{\text{GAN}}$ to guide the model in completing basic reconstruction tasks (Isola et al., 2016a; Ginosar et al., 2019). Additionally, we employ a KL divergence loss $L_{\text{KL}}$ to ensure that the latent code ${\bm{h}}$ closely approximates a standard Gaussian distribution.
@@
$\hat{{\bm{x}}}_{1:N}=\mathcal{D}(\text{DM}(\mathcal{E}({\bm{x}}_{\text{src}}),{\bm{y}}_{1:N},\bm{\rho}_{1:N}))$ (10)
@@ 3.6 Inference
Our inference process has four steps: 1) Extract the audio embedding ${\bm{a}}_{1:N}$. 2) Use the source image ${\bm{x}}_{\text{src}}$ to extract the initial pose/blink state $\bm{\rho}_{\text{src}}$ for PBNet. Along with ${\bm{a}}_{1:N}$ and a latent space vector ${\bm{h}}_{1:N}$ sampled from a standard Gaussian distribution, PBNet generates the pose/blink sequences $\hat{\bm{\rho}}_{1:N}$. 3) Input ${\bm{x}}_{\text{src}}$, ${\bm{a}}_{1:N}$, and $\hat{\bm{\rho}}_{1:N}$ into the A2V-FDM, which generates the motion representation sequences $\hat{{\bm{z}}}^{m}_{1:N}$. 4) Finally, decode the video sequence $\hat{{\bm{x}}}_{1:N}$ from ${\bm{x}}_{\text{src}}$ and $\hat{{\bm{z}}}^{m}_{1:N}$. Both PBNet and A2V-FDM generate sequences of dynamic lengths in a single pass, depending on input audio length.
@@ 4.1 Setup
Implementation details. The architecture of the encoder and decoder in our model aligns with the design proposed by Johnson et al. (2016), while the flow predictor is implemented based on MRAA (Siarohin et al., 2021). The PBNet model is trained using pose and blink movement sequences of 200 frames. During the inference phase of the PBNet model, a local attention mechanism with a window size of 400 is employed. For the inference phase of the A2V-FDM model, local attention with a window size of 80 is applied. In our evaluation, the length of one-time inference for CREMA is dynamic and depends on the ground truth, while for HDTF, it is fixed at 200 frames for better comparison.
Evaluation metrics. We evaluate the performance of our method using various metrics. Specifically, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) to assess the image quality. We utilize the $\text{FVD}_{16}$ and $\text{FVD}_{32}$ scores, which calculate the Fréchet Video Distance (FVD) based on window sizes of 16 and 32 frames, respectively, to evaluate video quality across different temporal scales. Furthermore, we assess the perception loss of lip shape using the confidence score ($\text{LSE}_{\text{C}}$) and distance score ($\text{LSE}_{\text{D}}$) (Chung & Zisserman, 2017). To evaluate the preservation of speaker identity, we use ArcFace (Deng et al., 2019) to extract features from both the ground truth image and the generated image, and use the cosine similarity (CSIM) between the two as the evaluation metric. Moreover, we employ the Beat Align Score (BAS) (Siyao et al., 2022) to evaluate the synchronization of head motion and audio, and calculate the number of blinks per second (blink/s) to assess the liveliness of the eyes. To better illustrate the error accumulation from a quantitative perspective, we design a metric to quantify the severity of error accumulation in the image space inspired by Bian et al. (2022) in the sequential generation task: Degradation Rate (DR), defined as $\text{DR}=\frac{\text{FID}_{\text{ed}}}{\text{FID}_{\text{st}}}-1$. DR is related to the ratio between the FID of the last $n$ frames $\text{FID}_{\text{ed}}$ and the first $n$ frames $\text{FID}_{\text{st}}$. The motivation for proposing this metric is that when error accumulation occurs, the quality of the generated data at the end of the sequence significantly deteriorates compared to the beginning. A larger DR indicates more severe error accumulation. In our experiments, we set $n$ to 25 frames (1s) and 50 frames (2s), denoted as $\text{DR}_{25}$ and $\text{DR}_{50}$, respectively.
@@ 4.2 Overall Comparison
We compared our method with several state-of-the-art methods: Wav2Lip (Prajwal et al., 2020), Audio2Head (Wang et al., 2021), SadTalker (Zhang et al., 2023), Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), Hallo (Xu et al., 2024b), and EchoMimic (Chen et al., 2024). The quantitative experimental results are shown in Table 1. Our method achieves promising performance in FID, $\text{FVD}_{16}$, $\text{FVD}_{32}$, CSIM, BAS, and Blink/s metrics for the CREMA and HDTF datasets, as shown in Table 1. It also attains nearly best scores for $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$. It is important to note that EchoMimic and DreamTalk are applicable only to face-aligned scenarios, whereas other methods do not require pre-cropping. Additionally, Hallo and EchoMimic are built upon pre-trained Stable Diffusion models (Rombach et al., 2022), inheriting substantial visual generation capabilities. Despite this, our method still achieves comparable or even better performance on the HDTF dataset. Furthermore, our method also outperforms Hallo and EchoMimic in terms of generation speed. In Appendix A.4.3, results indicate that our method achieves the fastest or near-fastest generation speed and requires significantly less time than Diffused Heads, Hallo, and EchoMimic.
Table 1: Quantitative comparison with several state-of-the-art methods methods on HDTF (Zhang et al., 2021) and CREMA (Cao et al., 2014) datasets. We use bold to indicate the best score and underline to represent the second-best score. * Wav2Lip generated videos that only contain lip motions, while the rest remain still images. For the sake of rigor, consider it a reference for the quality of lip motion and we will not include Wav2Lip in the ranking. $\uparrow$” indicates better performance with higher values, while $\downarrow$” indicates better performance with lower values. For both BAS and Blink/s, we consider performance to be better when they are closer to the ground truth.
@@
Table 2: Comparison with other generation strategies. The semi-autoregressive (SAR) generation strategy is similar to He et al. (2023). The two temporal resolution (TTR) generation method is mentioned in Harvey et al. (2022). For the DR metric, we consider performance to be better as it approaches zero.
@@
SAR 11.42 13.00 120.33 210.52 4.34 8.29 0.307 0.021
TTR 19.25 9.77 95.42 137.14 4.87 8.68 -0.028 -0.005
Ours 7.32 5.77 56.33 75.82 5.77 8.14 0.044 0.031
Table 3: The experiment of extrapolation evaluation. "Inference length" refers to the number of frames generated in a single inference process.
@@
Table 4: Ablation study on TCL and PBNet. The "GT PB" refers to whether to use ground truth pose/blink signal.
@@
For the qualitative experiment, as shown in Figure 2, we visualize the generation results for each baseline on different datasets. Our method evidently achieves the visual quality most similar to the ground truth, showcasing the most realistic and vivid visual effects. Compared to our method, SadTalker relies on the 3DMM prior, which limits its animation capability to the facial region only, resulting in significant artifacts when merging the face with the static torso below the neck. Additionally, SadTalker exhibits unnatural head pose movements and gaze direction, partially due to limited temporal modeling ability. Audio2Head fails to preserve the speaker’s identity during generation. For the HDTF dataset, the Diffused Heads method collapsed due to the error accumulation. DreamTalk is only applicable to face-aligned scenarios, thus requiring the source image to be cropped. Hallo leads to serious error accumulation in the CREMA dataset, resulting in abnormal color patches in the background, indicating a robustness defect.
@@
We compared our non-autoregressive generation strategy with two regular video generation strategies: 1) semi-autoregressive (SAR) generation similar to He et al. (2023), and 2) two-temporal resolution (TTR), which trains two models with different temporal resolutions (Harvey et al., 2022). The time cost represents the time required to generate an 8-second talking head video. The models were evaluated on the CREMA dataset, and the results are shown in Table 2. According to the results, our non-autoregressive method produces videos with the highest overall quality and fastest speed. In addition, the SAR method results in a very high DR, which is at least one order of magnitude higher than TTR and NAR. This is because autoregressive methods suffer from degradation issues during iterative generation. Although TTR also successfully alleviates the degradation issue, it compromises generation speed and overall quality. In summary, our non-autoregressive method addresses the degradation problem while preserving fast generation speed and high overall quality.